### v3

# CCDM: Continuous Conditional Diffusion Models for Image Generation

Abstract
Continuous Conditional Generative Modeling (CCGM) estimates high-dimensional data distributions, such as images, conditioned on scalar continuous variables (aka regression labels). While Continuous Conditional Generative Adversarial Networks (CcGANs) were designed for this task, their instability during adversarial learning often leads to suboptimal results. Conditional Diffusion Models (CDMs) offer a promising alternative, generating more realistic images, but their diffusion processes, label conditioning, and model fitting procedures are either not optimized for or incompatible with CCGM, making it difficult to integrate CcGANs’ vicinal approach. To address these issues, we introduce Continuous Conditional Diffusion Models (CCDMs), the first CDM specifically tailored for CCGM. CCDMs address existing limitations with specially designed conditional diffusion processes, a novel hard vicinal image denoising loss, a customized label embedding method, and efficient conditional sampling procedures. Through comprehensive experiments on four datasets with resolutions ranging from $64\times 64$ to $192\times 192$, we demonstrate that CCDMs outperform state-of-the-art CCGM models, establishing a new benchmark. Ablation studies further validate the model design and implementation, highlighting that some widely used CDM implementations are ineffective for the CCGM task. Our code is publicly available at https://github.com/UBCDingXin/CCDM.
Index Terms:
Continuous conditional generative modeling, conditional diffusion models, continuous scalar conditions.
I Introduction
Continuous Conditional Generative Modeling (CCGM), as depicted in Fig. 1, aims to model the probability distribution of images conditioned on scalar continuous variables. These variables, commonly referred to as regression labels, can encompass a diverse range of values, including angles, ages, counts, temperatures, coordinates, and more. However, CCGM faces significant challenges, particularly due to the scarcity of training samples for specific regression labels and the lack of an effective label input mechanism. These limitations adversely impact the performance of conventional Conditional Generative Adversarial Networks (cGANs) and Conditional Diffusion Models (CDMs), as demonstrated in [ding2023ccgan, ding2024turning].
Figure 1: Illustration of the CCGM task with sample images from the UTKFace and Steering Angle datasets.
As a recent advancement, Ding et al. introduced the first feasible model to the CCGM task, termed Continuous Conditional Generative Adversarial Networks (CcGANs) [ding2023ccgan]. CcGANs address the challenges of the CCGM task by developing novel vicinal discriminator losses and label input mechanisms. Consequently, CcGANs have been extensively applied across various domains requiring precise control over generative modeling of high-dimensional data. These applications comprise engineering inverse design [heyrani2021pcdgan, zhao2024ccdpm], data augmentation for hyperspectral imaging [zhu2023data], remote sensing image processing [giry2022sar], model compression [ding2023distilling], controllable point cloud generation [triess2022point], and more. Despite their success, CcGANs still face crucial limitations [ding2024turning], e.g., occasionally producing low-quality images marked by poor visual fidelity and label inconsistency—instances where generated images deviate from the specified conditions.
Diffusion models now lead image generation, consistently outperforming GANs in quality [dhariwal2021diffusion, ho2021classifier, rombach2022high, peebles2023scalable, esser2024scaling, labs2025flux1kontextflowmatching], making their adaptation to CCGM a promising path to surpass CcGANs [ding2023ccgan, ding2024turning]. However, existing Conditional Diffusion Models (CDMs) are designed for discrete conditions like class labels [dhariwal2021diffusion, ho2021classifier] or text descriptions [rombach2022high, peebles2023scalable, esser2024scaling, li2024g, qu2024discriminative, labs2025flux1kontextflowmatching], making them unsuitable for regression labels in CCGM due to numerous limitations: their diffusion processes lack an explicit condition incorporation mechanism, leading to suboptimal performance, particularly for label inconsistency and poor visual fidelity; their label input mechanisms require modification for regression labels; the commonly used noise prediction loss and hyperparameter selection schemes are largely ineffective in CCGM; and the soft vicinal loss optimal for CcGANs does not translate well to CDMs. Consequently, integrating CDMs with CcGAN techniques faces significant challenges, underscoring an urgent need for developing novel diffusion processes, loss functions, label input mechanisms, and training/sampling strategies specially tailored to CCGM.
Motivated by the aforementioned challenges, we introduce Continuous Conditional Diffusion Models (CCDMs) in this paper—a novel framework designed to overcome the crucial limitations of existing CDMs and offer a comprehensive guideline for their application in the CCGM task. Our key contributions are summarized as follows:
  • To improve control over generated images and enhance overall generation quality, we propose novel conditional diffusion processes that explicitly integrate conditional information through $y$-dependent state transitions, as detailed in Section III-A.
  • To address the data sparsity issue in CCGM, we derive the Hard Vicinal Negative Log-Likelihood (HV-NLL), inspired by the vicinal technique from CcGANs [ding2023ccgan]. Building on HV-NLL and the modified diffusion, we introduce the Hard Vicinal Image Denoising Loss (HVIDL) in Section III-C, a novel training loss that incorporates a hard vicinal weight and a $y$-dependent covariance matrix. We also argue that the widely used noise prediction loss [ho2020denoising] for diffusion models and the soft vicinal loss, while optimal for CcGANs, are unsuitable for CCGM—a claim supported by ablation studies in Section LABEL:sec:ablation.
  • In Section III-D, we employ a multi-step sampling algorithm optimized for CCGM, with configurations distinct from class-conditional or text-to-image generation. In Section III-E, we introduce DMD2-M, a one-step sampler inspired by DMD2 [yin2024improved], enhanced with advanced GAN architectures and training techniques to adapt it for the CCGM task. DMD2-M significantly accelerates sampling, with a slight trade-off in image quality.
  • Building on the comprehensive experiments in Sections IV and LABEL:sec:ablation, we demonstrate that CCDM achieves state-of-the-art performance. Additionally, we offer a thorough implementation guideline for CCDM.
It is worth mentioning that the proposed CCDM substantially differs from a recent work CcDPM [zhao2024ccdpm] in four aspects:
  • CcDPM focuses on non-image data, whereas high-dimensional image generation is significantly more challenging than non-image ones. Also, CcDPM was designed for multi-dimensional labels, lacking an efficient mechanism for scalar regression labels, whereas our CCDMs are specifically developed for image data with scalar continuous labels, making CCDM the first CDM tailored for the CCGM task.
  • CcDPM employs a noise prediction loss weighted by soft vicinity, while CCDM proposes a novel image denoising loss weighted by hard vicinity, which our ablation study in Section LABEL:sec:ablation demonstrates significantly better performance in CCGM tasks.
  • CcDPM’s diffusion processes are not directly tied to regression labels, whereas CCDM explicitly integrates regression labels into both forward and reverse processes, enhancing image quality.
  • CCDM provides a mathematically rigorous derivation of its modified diffusion processes and training loss in Appendix LABEL:sec:supp_detailed_derivations, which is absent in CcDPM. These distinctions underscore CCDM’s unique effectiveness in image generation with scalar continuous labels.
II Related Work
II-A Continuous Conditional GANs
CcGANs, introduced by Ding et al. [ding2021ccgan, ding2023ccgan], represent the pioneering approach to the Continuous Conditional Generative Modeling (CCGM) task. As shown in Fig. 1, mathematically, CcGANs are devised to estimate the probability density function $p(\bm{x}^{0}|y)$, characterizing the underlying conditional data distribution, where $y$ represents regression labels. In addressing concerns regarding the potential data insufficiency at $y$, Ding et al. [ding2021ccgan, ding2023ccgan] posited an assumption wherein minor perturbations to $y$ result in negligible alterations to $p(\bm{x}^{0}|y)$. Building upon this assumption, they formulated the Hard and Soft Vicinal Discriminator Losses (HVDL and SVDL) alongside a novel generator loss, aimed at enhancing the stability of GAN training. Furthermore, to overcome the challenge of encoding regression labels, Ding et al. [ding2021ccgan, ding2023ccgan] proposed both a Naïve Label Input (NLI) mechanism and an Improved Label Input (ILI) mechanism to effectively integrate $y$ into CCGM models. The ILI approach employs a label embedding network $\phi$ comprising a 5-layer perceptron to map the scalar $y$ to a $128\times 1$ vector $\bm{h}^{s}_{y}$, i.e.,$\bm{h}^{s}_{y}=\phi(y)$. This vector, $\bm{h}^{s}_{y}$, is then fed into the generator and discriminator networks of CcGANs to control the generative modeling process. The efficacy of CcGANs has been demonstrated by Ding et al. [ding2021ccgan, ding2023ccgan] and other researchers [heyrani2021pcdgan, giry2022sar, zhu2023data] across diverse datasets, showcasing their robustness and versatility in various applications. However, as pointed out in [ding2024turning], the adversarial training mechanism is vulnerable to data insufficiency, and CcGANs may still generate subpar fake images. Therefore, in this paper, we propose to transition away from the conventional adversarial learning paradigm and instead adopt conditional diffusion processes for CCGM, meanwhile improving the vicinal training and label input mechanisms.
II-B Diffusion Models
As an alternative to GANs, diffusion models offer a viable approach for generative modeling and are widely used in computer vision tasks [croitoru2023diffusion, 10480591, 10261222, 10321681]. A notable example is Denoising Diffusion Probabilistic Models (DDPMs) [ho2020denoising], which consist of forward and reverse diffusion processes modeled as Markov chains. The forward diffusion process gradually transforms a real image $\bm{x}^{0}\sim p(\bm{x}^{0})$ into pure Gaussian noise $\bm{x}^{T}\sim\mathcal{N}(\bm{0},\bm{I})$ through $T$ steps, defined by transitions $q(\bm{x}^{t}|\bm{x}^{t-1})$, where $T$ is sufficiently large. The reverse process, parameterized by $\bm{\theta}$, reconstructs $\bm{x}^{0}$ from $\bm{x}^{T}$ through transitions $p_{\bm{\theta}}(\bm{x}^{t-1}|\bm{x}^{t})$. DDPMs are trained using a noise prediction (aka $\bm{\epsilon}$-prediction) loss defined as follows:
$\mathcal{L}_{simple}(\bm{\theta})=\mathbb{E}_{\bm{\epsilon},\bm{x}^{t}}\|\hat{\bm{\epsilon}}_{\bm{\theta}}(\bm{x}^{t},t)-\bm{\epsilon}\|^{2}_{2}$ (1)
where $\hat{\bm{\epsilon}}_{\bm{\theta}}(\bm{x}^{t},t)$ is modeled by a U-Net [ronneberger2015u]. This U-Net is trained to predict the ground truth sampled Gaussian $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. Although DDPMs can generate high-quality data, they suffer from long sampling times. To mitigate this, Song et al. proposed Denoising Diffusion Implicit Models (DDIMs) [song2021denoising], which introduce a deterministic yet shorter reverse process for high-quality sample generation with significantly fewer steps.
However, DDIM still typically requires hundreds of steps for sample generation, which remains time-consuming. To enable few-step or even one-step sampling, distillation-based methods are employed, with Improved Distribution Matching Distillation (DMD2) [yin2024improved] representing a state-of-the-art approach. DMD2 comprises a generator, a real score function, a fake score function, and a discriminator, all modeled with convolutional neural networks. The real score function is a fixed pre-trained U-Net from the diffusion model, while the generator and fake score function, also based on this U-Net, are trainable. The discriminator’s feature extraction network shares weights with the encoder of the fake score function and is followed by fully connected layers that output the probability of an input image being real. The training process for DMD2 alternates between two steps: first, optimizing the generator using a distribution matching loss combined with the vanilla GAN loss [goodfellow2014generative]; second, training the fake score function and the discriminator with an image denoising loss and the vanilla GAN loss. Once training is complete, the generator is used for image generation. Notably, DMD2 multiplies the GAN losses for both the discriminator and generator by two positive weights, $w_{D}$ and $w_{G}$, which are treated as hyperparameters.
Furthermore, unlike DDPMs, which aim to estimate the marginal data distribution $p(\bm{x}^{0})$, Ho et al. [ho2021classifier] introduced the Classifier-Free Guidance (CFG) model to estimate the conditional distribution $p(\bm{x}^{0}|y)$. As a CDM, CFG integrates class labels as conditions and employs the following equation for sampling [luo2022understanding]:
$\triangledown\log p(\bm{x}^{t}|y)=(1-\gamma)\triangledown\log p(\bm{x}^{t})+\gamma\triangledown\log p(\bm{x}^{t}|y)$ (2)
where $\triangledown\log p(\bm{x}^{t}|y)$ is an abbreviation of the score function $\triangledown_{\bm{x}^{t}}\log p(\bm{x}^{t}|y)$, $y$ represents class labels, and $\gamma$ is a positive hyper-parameter controlling the model’s consideration of conditioning information during the sampling procedure.
II-C Text-to-Image Generation
Text-to-image (T2I) generation similarly involves conditional distribution estimation, driving advances in modern diffusion models [rombach2022high, peebles2023scalable, esser2024scaling, labs2025flux1kontextflowmatching]. However, as noted in [ding2023ccgan, ding2024turning] and demonstrated in our experimental results (Section IV), current T2I architectures like Stable Diffusion [rombach2022high] exhibit critical limitations in CCGM tasks. First, despite their ability to produce realistic images from prompts, these models inherently lack the architectural capacity to address the challenge of scarce or even zero training data for specific regression labels. Second, their dependence on large text encoders like CLIP [radford2021learning] for prompt embedding—significantly more complex than CcGAN’s compact 5-layer perceptron $\phi(y)$—introduces substantial and unnecessary memory overhead in CCGM applications.
Recent advances in text-to-image alignment [qu2024discriminative, li2024g] have explored quantitative prompt conditioning, though primarily focusing on discrete counting tasks. These approaches, while making progress in their specific domains, prove inadequate for addressing CCGM requirements. A key limitation lies in their inability to properly model the conditional distribution $p(\bm{x}|y)$ for regression labels or text containing such labels. The current frameworks are typically constrained to processing discrete numerical values (e.g., “Two girls sitting on the grass”) rather than continuous regression labels such as angular measurements. More significantly, these methods inherit the same fundamental limitations as general T2I diffusion models (including Stable Diffusion) when applied to CCGM tasks, failing to overcome the core challenges of conditional generation in this domain.
Consequently, quantitative prompt-conditioned image generation remains a largely unexplored area of research.
III Methodology
In this section, we introduce CCDM, a novel approach specially designed to estimate the conditional distribution $p(\bm{x}^{0}|y)$ based on $N$ image-label pairs $\{(\bm{x}^{0}_{i},y_{i})\}_{i=1}^{N}$. CCDM introduces a new $y$-dependent diffusion processes, proposes a novel hard vicinal image denoising loss for training, adapts the CcGAN’s label input mechanism to diffusion models, and proposes two efficient sampling algorithms. The overall workflow is depicted in Fig. 2, where novel or customized components are highlighted in red.
Figure 2: The overall workflow of CCDM. Our method introduces $y$-dependent conditional diffusion processes and a novel hard vicinal image denoising loss (HVIDL) for U-Net training. To integrate the regression label $y$ into CCDM, we use two types of embeddings: a short embedding $\bm{h}_{y}^{s}$, encoded by CcGAN’s embedding network for U-Net injection, and a long embedding $\bm{h}_{y}^{l}$, encoded by the proposed covariance embedding network for forward diffusion and HVIDL. For inference, we employ a multi-step sampling algorithm combining CFG [ho2021classifier] and DDIM [song2021denoising], specially optimized for CCGM. We further introduce DMD2-M, an adaptation of DMD2 [yin2024improved] for one-step CCGM sampling with minimal quality trade-off.
III-A Conditional Diffusion Processes
The proposed CCDM is a conditional diffusion model that comprises two key components: a forward diffusion process and a reverse diffusion process. In both processes, the regression label $y$ is explicitly incorporated as a conditioning factor.
III-A1 Forward Diffusion Process
Given a random sample $\bm{x}^{0}$ drawn from the actual conditional distribution $\bm{x}^{0}\sim p(\bm{x}^{0}|y)$, we define a forward diffusion process by incrementally adding a small amount of $y$-dependent Gaussian noise over $T$ steps. This process generates a Markov chain comprising a sequence of noisy samples $\bm{x}^{1},\ldots,\bm{x}^{T}$. All noisy samples in this sequence retain the same regression labels $y$ and dimensionality as $\bm{x}^{0}$. The state transition for this Markov chain is defined as the following $y$-dependent Gaussian model:
$q(\bm{x}^{t}|\bm{x}^{t-1},y)=\mathcal{N}(\bm{x}^{t};\sqrt{1-\beta_{t}}\bm{x}^{t-1},\beta_{t}\bm{H}_{y})$ (3)
where $\bm{H}_{y}=\text{diag}(\exp(-\bm{h}^{l}_{y}))$ is a diagonal covariance matrix, and $\bm{h}^{l}_{y}$ is a high-dimensional embedding of label $y$ with the same length as flattened $\bm{x}^{0}$. We train a covariance embedding network, denoted as $\phi^{\prime}$, comprising five fully-connected layers, to transform the scalar $y$ to the vector $\bm{h}^{l}_{y}$, expressed as $\bm{h}^{l}_{y}=\phi^{\prime}(y)$. Further details of $\phi^{\prime}$ can be found in Section III-B and Appendix LABEL:sec:supp_cov_label_embedding. Additionally, the parameters $0\leq\beta_{1}<...<\beta_{T}\leq 1$ are fixed to constants using the cosine schedule [nichol2021improved]. Let $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, we can sample $\bm{x}^{t}$ at any arbitrary time step $t$ from
$q(\bm{x}^{t}|\bm{x}^{0},y)=\mathcal{N}(\bm{x}^{t};\sqrt{\bar{\alpha}_{t}}\bm{x}^{0},(1-\bar{\alpha}_{t})\bm{H}_{y})$ (4)
Therefore, we have
$\bm{x}^{t}=\sqrt{\bar{\alpha}_{t}}\bm{x}^{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}\quad\text{where}\quad\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{H}_{y})$ (5)
III-A2 Reverse Diffusion Process
In the reverse diffusion process, we aim to leverage the condition $y$ to guide the reconstruction of $\bm{x}^{0}$ from the Gaussian noise $\bm{x}^{T}\sim\mathcal{N}(\bm{0},\bm{H}_{y})$. Unfortunately, estimating $q(\bm{x}^{t-1}|\bm{x}^{t},y)$ is usually nontrivial. Nevertheless, this reversed conditional distribution is tractable when conditioned on $\bm{x}^{0}$:
$q(\bm{x}^{t-1}|\bm{x}^{t},\bm{x}^{0},y)=\mathcal{N}(\bm{x}^{t-1};\bm{\mu}_{q}(\bm{x}^{t},\bm{x}^{0},y),\bm{\Sigma}_{q}(t,y)).$ (6)
The conditional distribution above is known as the ground truth denoising transition. Specifically,
$\bm{\mu}_{q}(\bm{x}^{t},\bm{x}^{0},y)=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\bm{x}^{t}+\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})\bm{x}^{0}}{1-\bar{\alpha}_{t}}$ (7)
$\bm{\Sigma}_{q}(t,y)=\frac{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\bm{H}_{y}\triangleq\sigma_{q}^{2}(t)\bm{H}_{y}.$ (8)
The derivation of Eq. (6) can be found in Appendix LABEL:sec:supp_derivation_reverse_process.
When generating new data points, $\bm{x}^{0}$ is inherently unknown. Therefore, we train a parametric model to approximate $q(\bm{x}^{t-1}|\bm{x}^{t},\bm{x}^{0},y)$ using the following form:
$p_{\bm{\theta}}(\bm{x}^{t-1}|\bm{x}^{t},y)=\mathcal{N}(\bm{x}^{t-1};\bm{\mu}_{\bm{\theta}}(\bm{x}^{t},t,y),\bm{\Sigma}_{q}(t,y))$ (9)
where,
$\bm{\mu}_{\bm{\theta}}(\bm{x}^{t},t,y)$ (10)
$=$ $\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})\bm{x}^{t}+\sqrt{\bar{\alpha}_{t-1}(1-\alpha_{t})}\hat{\bm{x}}_{\bm{\theta}}^{0}(\bm{x}^{t},t,y)}{1-\bar{\alpha}_{t}}$ (11)
In Eq. (11), $\hat{\bm{x}}_{\bm{\theta}}^{0}(\bm{x}^{t},t,y)$ is parameterized by a denoising U-Net [ronneberger2015u, ho2020denoising], designed to predict $\bm{x}^{0}$ based on the noisy image $\bm{x}^{t}$, timestamp $t$, and condition $y$ associated with both $\bm{x}^{t}$ and $\bm{x}^{0}$, where $\bm{\theta}$ denotes the trainable parameters. Further details on this U-Net are provided in Appendix LABEL:sec:supp_unet_arch.
Remark 1.
For practical demonstration, we opt to use a DDPM-based formulation of diffusion processes, despite the existence of various alternative explanations of diffusion models, such as those grounded in Stochastic Differential Equations (SDEs) [song2019generative, song2021scorebased]. Furthermore, in contrast to DDPM, the diffusion processes presented above explicitly integrate conditional information $y$ through the $y$-dependent covariance matrix $\bm{H}_{y}$.
III-B Covariance Embedding Network
Although the denoising U-Net for DDPM [ronneberger2015u, ho2020denoising] supports the sinusoidal positional embedding [ho2020denoising] and Gaussian Fourier projection [song2021scorebased] for encoding time steps $t$, Table LABEL:tab:ab_effect_embedding shows that these methods are ineffective for embedding regression labels $y$. As illustrated in Fig. 2, $y$ comprises two types of embeddings: (1) a short embedding $\bm{h}_{y}^{s}\in\mathbb{R}^{128}$, which serves as input to the denoising U-Net, and (2) a long embedding $\bm{h}_{y}^{l}\in\mathbb{R}^{(C\times W\times H)}$, which modulates the diffusion processes and contributes to both CCDM’s training loss and sampling algorithms, where $C$, $W$, and $H$ denote the number of channels, width, and height of the input images, respectively. We utilize CcGAN’s embedding network $\phi(y)$ [ding2023ccgan] to convert the scalar $y$ into a 128-dimensional vector $\bm{h}^{s}_{y}$. However, $\phi(y)$ is unsuitable for computing $\bm{h}_{y}^{l}$ due to the significantly higher dimension of $\bm{h}_{y}^{l}$ compared to $\bm{h}_{y}^{s}$. To address this, we adapt CcGAN’s label embedding network to create the covariance embedding network, denoted as $\phi^{\prime}(y)$, to convert $y$ into $\bm{h}_{y}^{l}$. We employ a training scheme nearly identical to [ding2023ccgan] to train $\phi^{\prime}(y)$, with modifications to the architecture and a reduction in the training epochs of an auxiliary ResNet34 network to suit our requirements. For a more detailed explanation, please refer to Appendix LABEL:sec:supp_cov_label_embedding.
III-C Hard Vicinal Image Denoising Loss
The Negative Log-Likelihood (NLL) of CCDM along with its upper bound are defined as follows
$-\log\mathbb{E}_{y\sim p(y)}\left[\prod_{k=1}^{N^{y}}f_{\bm{\theta}}(\bm{x}_{k,y}^{0}|y)\right]$ (12)
$\leq$ $\mathbb{E}_{y\sim p(y)}\left[\sum_{i=1}^{N}\mathds{1}_{\{y_{i}=y\}}\left(-\log f_{\bm{\theta}}(\bm{x}_{i}^{0}|y)\right)\right]$ (13)
where $f_{\bm{\theta}}(\bm{x}^{0}|y)$ is the likelihood function for observed images with label $y$, $\bm{x}_{k,y}^{0}$ is the $k$-th image with label $y$, $N^{y}$ stands for the number of images with label $y$, $p(y)$ is the label distribution, and $\mathds{1}$ is the indicator function. Note that Eq. (13) suggests that estimating $p(\bm{x}^{0}|y)$ relies solely on images with label $y$, potentially suffering from the data insufficiency issue.
To address data sparsity, we incorporate the vicinal loss from CcGANs into the training of conditional diffusion models. Assume we use training images with labels in a hard vicinity of $y$ to estimate $p(\bm{x}^{0}|y)$, then we adjust Eq. (13) to derive the Hard Vicinal NLL (HV-NLL) as follows:
$\mathbb{E}_{y\sim p(y)}\left[\sum_{i=1}^{N}\mathds{1}_{\{|y_{i}-y|\leq\kappa\}}\left(-\log f_{\bm{\theta}}(\bm{x}_{i}^{0}|y)\right)\right]$ (14)
$\approx$ $\frac{C}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}\left\{\mathbb{E}_{\delta\sim\mathcal{N}(0,\sigma_{\delta}^{2})}\left[W_{h}\cdot(-\log f_{\bm{\theta}}(\bm{x}_{i}^{0}|y_{j}+\delta))\right]\right\}$ (15)
where $\sigma_{\delta}$ is the bandwidth for Kernel Density Estimation (KDE) of $p(y)$, $C$ is a constant, and
$W_{h}\triangleq\mathds{1}_{\{|y_{j}+\delta-y_{i}|\leq\kappa\}}.$ (16)
Although not recommended, the hard vicinal weight $W_{h}$ can be replaced with a soft vicinal weight as follows:
$W_{s}\triangleq e^{-\nu(y_{j}+\delta-y_{i})^{2}}.$ (17)
Eq. (15) is not analytically tractable, so we minimize its upper bound instead of directly optimizing it. After derivations and simplifications (see Appendix LABEL:sec:supp_detailed_derivation_loss), we obtain the Hard Vicinal Image Denoising Loss (HVIDL) for training the U-Net:
[…]
$\left[W_{h}\cdot(\hat{\bm{x}}_{\bm{\theta}}^{0}-\bm{x}^{0}_{i})^{\intercal}\bm{H}_{y_{j}+\delta}^{-1}(\hat{\bm{x}}_{\bm{\theta}}^{0}-\bm{x}^{0}_{i})\right]\right\}.$ (19)
Eq.(19) represents an image denoising (aka $\bm{x}^{0}$-prediction) loss that is weighted by both the hard vicinal weight $W_{h}$ and the inverse of the $y$-dependent covariance matrix $\bm{H}_{y_{j}+\delta}^{-1}$. We minimize this loss for model fitting, with the detailed training procedure outlined in Algorithm 1.
Remark 2.
The positive hyperparameters $\kappa$, $\nu$, and $\sigma_{\delta}$ can be determined using the rule of thumb in Remark 3 of [ding2023ccgan].
Remark 3.
[…]
Remark 4.
The ablation study in Section IV shows that the hard vicinal weight $W_{h}$ outperforms the soft one $W_{s}$. Thus, we recommend using the hard vicinal weight $W_{h}$ in applications.
Remark 5.
It can be shown that Eq. (19) can be transformed into a $\bm{\epsilon}$-prediction loss by utilizing the relationship between $\bm{x}^{0}$-prediction and $\bm{\epsilon}$-prediction [ho2020denoising, luo2022understanding]. However, when using the accelerated sampler DDIM [song2021denoising], this $\bm{\epsilon}$-prediction loss often leads to significant label inconsistency, as evidenced in Table LABEL:tab:ab_effect_objective.
III-D CFG and DDIM-based Multi-Step Sampling
To generate samples from the learned model, we adapt the CFG framework [dhariwal2021diffusion], originally designed for class-conditional sampling, to CCGM. Given a regression label $y$, we aim to generate corresponding fake images. Luo [luo2022understanding] demonstrated a linear relationship between $\bm{x}^{0}$ and $\triangledown\log p(\bm{x}^{t}|y)$ as follows:
$\bm{x}^{0}=\frac{\bm{x}^{t}+(1-\bar{\alpha}_{t})\triangledown\log p(\bm{x}^{t}|y)}{\sqrt{\bar{\alpha}_{t}}}$ (20)
We then modify Eq. (2) to reconstruct $\bm{x}^{0}$:
$\tilde{\bm{x}}^{0}=(1-\gamma)\cdot\underbrace{\hat{\bm{x}}_{\bm{\theta}}^{0}(\bm{x}^{t},t,\oslash)}_{\text{unconditional}}+\gamma\cdot\underbrace{\hat{\bm{x}}_{\bm{\theta}}^{0}(\bm{x}^{t},t,y)}_{\text{conditional}},$ (21)
where $\gamma$ (referred to as the conditional scale) is recommended as $\gamma\in[1.5,2]$, alongside $p_{\text{drop}}=0.1$, to balance image diversity and label consistency in the CCGM task. This contrasts with class-conditional or text-to-image scenarios, where typical settings are $\gamma>3$ and $p_{\text{drop}}\geq 0.2$ [dhariwal2021diffusion, balaji2022ediff].
For fast sampling, we use the DDIM sampler [song2021denoising] to generate new samples with $T^{\prime}\ll T$ time steps through the following deterministic procedure:
$\bm{x}^{t-1}=\sqrt{\bar{\alpha}_{t-1}}\tilde{\bm{x}}^{0}+\sqrt{1-\bar{\alpha}_{t-1}}\frac{\bm{x}^{t}-\sqrt{\bar{\alpha}_{t}}\tilde{\bm{x}}^{0}}{\sqrt{1-\bar{\alpha}_{t}}},$ (22)
where the initial $\bm{x}^{T^{\prime}}$ is randomly sampled from $\mathcal{N}(\bm{0},\bm{H}_{y})$ instead of a standard Gaussian $\mathcal{N}(\bm{0},\bm{I})$, making the sampling algorithm (Algorithm 2) different from [song2021denoising].
III-E Distillation-based One-Step Sampling
Even with DDIM, conditional sampling with CFG remains time-consuming, often taking hours or days to generate sufficient evaluation samples. To enhance efficiency, we attempted to distill the trained CCDM into a one-step generator using DMD2 [yin2024improved], but its model design and training approach proved ineffective for our CCGM task. For instance, when applied to RC-49 ($64\times 64$), we observed severe mode collapse, as shown in Fig. 3 and Table LABEL:tab:ablation_dmd2m. Furthermore, one-step sampling does not require a U-Net architecture, and sharing the feature extraction head between the discriminator and fake score function harms the discriminator’s performance, as these models need distinct feature representations. Additionally, DMD2’s reliance on vanilla GAN loss and outdated training techniques makes it unsuitable for CCGM tasks, which often face data insufficiency and imbalance. To address these issues, we propose DMD2-M, an enhanced version of DMD2 with the following improvements:
  • Network architecture: DMD2-M adopts the SNGAN [miyato2018spectral] architecture for both generator $G$ and discriminator $D$, using a 128-dimensional Gaussian noise for the generator instead of a high-dimensional noisy image. The discriminator is trained independently and does not share a feature extraction head with the fake score function model.
  • Loss function: DMD2-M significantly improves upon DMD2 by adopting the hinge GAN loss [miyato2018spectral] in place of the vanilla GAN loss. Moreover, it incorporates the hard vicinal weight (Eq. (16)) into all training losses to help mitigate mode collapse in data-limited scenarios. Additionally, our proposed $y$-dependent covariance matrix $\bm{H}_{y}$ is integrated into the forward diffusion process for the computation of training losses.
  • Training technique: DMD2-M integrates DiffAugment [zhao2020differentiable] into the distillation process—a highly effective and user-friendly technique that stabilizes adversarial training and improves performance, especially in situations with limited training data.
Remark 6.
By default, CCDM utilizes DDIM as its sampler to ensure optimal image quality unless stated otherwise. In contrast, DMD2-M is tailored for scenarios demanding rapid sampling, where a slight compromise in quality and extended training time are acceptable.
Figure 3: DMD2 [yin2024improved] suffers from severe mode collapse on RC-49 ($64\times 64$), demonstrating its inapplicability to the CCGM task.
Data: $N$ real image-label pairs $\Omega=\{(\bm{x}^{0}_{i},y_{i})\}_{i=1}^{N}$, $\tilde{N}$ distinct labels $\Upsilon=\{y_{[i]}\}_{i=1}^{\tilde{N}}$ arranged in ascending order from the training dataset;
Input: The pre-trained CcGAN’s embedding network $\phi(y)$ [ding2023ccgan]; The pre-trained covariance embedding network $\phi^{\prime}(y)$; Preset hard vicinity radius $\kappa$, KDE bandwidth $\sigma_{\delta}$, time steps $T$, condition drop probability $p_{\text{drop}}$, number of iterations $K$, and batch size $m$;
Result: A trained U-Net $\hat{\bm{x}}_{\bm{\theta}}^{0}$
1 for $k=1$ to $K$ do
2       Draw $m$ regression labels, denoted by $Y$, with replacement from $\Upsilon$;
3       Create a batch of target labels $Y^{\delta}=\{y_{i}+\delta|y_{i}\in Y,\delta\sim\mathcal{N}(0,\sigma_{\delta}^{2}),i=1,\dots,m\}$ ($\hat{\bm{x}}_{\bm{\theta}}^{0}$ is conditional on these labels) ;
4       Initialize $\Omega^{y_{i}+\delta}=\emptyset$ ;
5       for $i=1$ to $m$ do
6             Randomly choose an image-label pair $(\bm{x}^{0},y)\in\Omega$ satisfying $|y-y_{i}-\delta|\leq\kappa$ where $y_{i}+\delta\in Y^{\delta}$;
7             Draw $u\sim U(0,1)$;
8             if $u\geq p_{\text{drop}}$ then
9                   Update $\Omega^{y_{i}+\delta}=\Omega^{y_{i}+\delta}\cup(\bm{x}^{0},y_{i}+\delta)$;
10
11             else
12                   Update $\Omega^{y_{i}+\delta}=\Omega^{y_{i}+\delta}\cup(\bm{x}^{0},\emptyset)$;
13
14             end if
15
16       end for
17      Convert regression labels in $\Omega^{y_{i}+\delta}$ into their multi-dimensional representations $\bm{h}_{y_{i}+\delta}^{s}$ and $\bm{h}_{y_{i}+\delta}^{l}$ by applying the functions $\phi(y)$ and $\phi^{\prime}(y)$, respectively ;
18       Compute the $y$-dependent covariance matrix $\bm{H}_{y_{i}+\delta}$ via $\bm{H}_{y_{i}+\delta}=\text{diag}(\exp(-\bm{h}^{l}_{y_{i}+\delta}))$ ;
19       Update $\hat{\bm{x}}_{\bm{\theta}}^{0}$ using samples in the set $\Omega^{y_{i}+\delta}$ via gradient-based optimizers according to Eq. (19);
20
21 end for
Algorithm 1 Algorithm for Training CCDM
Input: The pre-trained U-Net $\hat{\bm{x}}_{\bm{\theta}}^{0}$; The pre-trained CcGAN’s embedding network $\phi(y)$ [ding2023ccgan]; The pre-trained covariance embedding network $\phi^{\prime}(y)$; The target regression label $y$; Preset hyperparameters, including the time steps $T^{\prime}$ for sampling, the conditional scale $\gamma$, and the variance schedule $\beta_{i}$’s;
Result: A fake image $\bm{x}^{0}_{g}$ with label $y$
1 Convert the regression label $y$ into its multi-dimensional representations $\bm{h}_{y}^{s}$ and $\bm{h}_{y}^{l}$ by applying the functions $\phi(y)$ and $\phi^{\prime}(y)$, respectively ;
2 Compute the $y$-dependent covariance matrix $\bm{H}_{y}$ via $\bm{H}_{y}=\text{diag}(\exp(-\bm{h}^{l}_{y}))$ ;
3 $t\leftarrow T^{\prime}$, $\bm{x}^{T^{\prime}}_{g}\sim\mathcal{N}(\bm{0},\bm{H}_{y})$ ;
4 while $t>0$ do
5       Compute $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$;
6       Compute $\tilde{\bm{x}}^{0}=(1-\gamma)\cdot\hat{\bm{x}}_{\bm{\theta}}^{0}(\bm{x}^{t}_{g},t,\oslash)+\gamma\cdot\hat{\bm{x}}_{\bm{\theta}}^{0}(\bm{x}^{t}_{g},t,y)$;
7       Update $\bm{x}^{t-1}_{g}=\sqrt{\bar{\alpha}_{t-1}}\tilde{\bm{x}}^{0}+\sqrt{1-\bar{\alpha}_{t-1}}\frac{\bm{x}^{t}_{g}-\sqrt{\bar{\alpha}_{t}}\tilde{\bm{x}}^{0}}{\sqrt{1-\bar{\alpha}_{t}}}$;
8       Update $t\leftarrow t-1$;
9 end while
Algorithm 2 Multi-Step Sampling Algorithm for CCDM Utilizing CFG and DDIM
IV Experiments
IV-A Experimental Setup
Datasets. We evaluate the effectiveness of CCDM across four datasets with varying resolutions ranging from $64\times 64$ to $192\times 192$. These datasets comprise RC-49 [ding2023ccgan], UTKFace [utkface], Steering Angle [steeringangle], and Cell-200 [ding2023ccgan]. Brief introductions to these datasets are presented below:
  • RC-49: The RC-49 dataset consists of 44,051 RGB images of 49 chair types at $64\times 64$ resolution. Each chair type is labeled with 899 images at yaw rotation angles ranging from $0.1^{\circ}$ to $89.9^{\circ}$ in $0.1^{\circ}$ increments. For training, we select angles with odd last digits and randomly sample 25 images per angle, resulting in a training set of 11,250 images corresponding to 450 distinct yaw angles.
  • UTKFace: The UTKFace dataset comprises 14,760 RGB human face images with ages as regression labels, spanning from 1 to 60 years [ding2021ccgan]. It features 50 to 1051 images per age, all used for training. The dataset is available in three resolutions: $64\times 64$, $128\times 128$, and $192\times 192$.
  • Steering Angle: The Steering Angle dataset, derived from an autonomous driving dataset [steeringangle], is preprocessed by Ding et al. [ding2023ccgan] for our experiments. It consists of 12,271 RGB images captured by a dashboard-mounted camera, available in resolutions of $64\times 64$ and $128\times 128$. These images are labeled with 1,774 unique steering wheel rotation angles, ranging from $-80.0^{\circ}$ to $80.0^{\circ}$.
  • Cell-200: The Cell-200 dataset, created by Ding et al. [ding2023ccgan], contains 200,000 synthetic grayscale microscopic images at a $64\times 64$ resolution. Each image contains a variable number of cells, ranging from 1 to 200, with 1,000 images available for each cell count. For training, we select a subset with only images containing an odd number of cells, choosing 10 images per cell count, resulting in 1,000 training images. For evaluation, we use all 200,000 samples from the dataset.
Compared Methods. In accordance with the setups described in [ding2024turning], we choose the following contemporary conditional generative models for comparison:
  • Two class-conditional GANs: We compare against ReACGAN [kang2021rebooting] and ADCGAN [hou2022conditional], both leveraging state-of-the-art architectures and training techniques with discretized regression labels.
  • Two class-conditional diffusion models: For comparison, we select two classic CDMs, ADM-G [dhariwal2021diffusion] and Classifier-Free Guidance (CFG)[ho2021classifier], both conditioned on discretized regression labels. We exclude more advanced CDMs like Stable Diffusion (SD)[rombach2022high] and DiT [peebles2023scalable] for the following reasons: This study serves as a proof of concept, focusing on demonstrating how our modifications to diffusion processes, conditioning mechanisms, and the introduction of a new vicinal loss enable effective application of CDMs to CCGM. These enhancements are applicable to state-of-the-art CDMs like SD and DiT but require substantial computational resources beyond our current capabilities. Therefore, we base CCDM on DDPM [ho2020denoising] and CFG [ho2021classifier] and ensure a fair comparison by evaluating it against ADM-G and CFG, which have similar model capacities.
  • A text-to-image diffusion model: We include Stable Diffusion v1.5 (SD v1.5) [rombach2022high] in our comparative evaluation on RC-49 and UTKFace, performing full fine-tuning of the official Hugging Face checkpoint (stable-diffusion-v1-5).
  • Two GAN-based CCGM models: We compare against state-of-the-art CcGAN (SVDL+ILI) [ding2023ccgan] and Dual-NDA [ding2024turning], a data augmentation strategy designed to enhance CcGAN’s visual quality and label consistency. Both methods are based on the CcGAN framework.
  • Two diffusion-based CCGM models: We include CcDPM [zhao2024ccdpm] and our proposed CCDM in the comparison. CcDPM, originally designed for non-image data, is adapted for the CCGM task by replacing its U-Net backbone and label embedding mechanism with our proposed approach. However, we retained CcDPM’s soft vicinity and $\bm{\epsilon}$-prediction loss.
For the $192\times 192$ experiment, we focus exclusively on SD v1.5, CcGAN (SVDL+ILI), Dual-NDA, CcDPM, and CCDM.
Implementation Setup. In our implementation of class-conditional models, we bin the regression labels of four datasets into various classes: 150 classes for RC-49, 60 classes for UTKFace, 221 classes for Steering Angle, and 100 classes for Cell-200. To facilitate comparisons, we leveraged pre-trained models provided by Ding et al.[ding2024turning] for experiments on UTKFace ($64\times 64$ and $128\times 128$) and Steering Angle ($64\times 64$ and $128\times 128$) datasets, rather than re-implementing all candidate methods (excluding CCDM). Additionally, Ding et al.[ding2023ccgan] offered pre-trained ReACGAN for RC-49 and CcGAN (SVDL+ILI) for UTKFace ($192\times 192$). For other candidate models, we re-implemented them for comparison. When implementing CcGAN (SVDL+ILI) and Dual-NDA on Cell-200, we used DCGAN [radford2015unsupervised] as the backbone and trained with the vanilla cGAN loss [mirza2014conditional]. For RC-49, we employed SAGAN [zhang2019self] as the backbone and trained with hinge loss [lim2017geometric] and DiffAugment [zhao2020differentiable]. For CCDM, we set $p_{\text{drop}}=0.1$ in Remark 3 and $\gamma=1.5$ in Eq. (21) for most experiments, except for UTKFace ($128\times 128$ and $192\times 192$), where we increased $\gamma$ to 2.0 for better label consistency. The total time step $T$ is fixed at 1000 during training, while the sampling time step $T^{\prime}$ varies: 250 for the $64\times 64$ experiments, and 150 for all other experiments. The hyperparameters $\kappa$ and $\sigma_{\delta}^{2}$ were set according to the rule of thumb in Remark 3 of [ding2023ccgan]. The training configurations of CcDPM were largely similar to those of CCDM, with the exception that CcDPM utilized $\bm{\epsilon}$-prediction loss and employed the soft vicinal weight. The one-step sampler, DMD2-M, is implemented only in the RC-49 ($64\times 64$) and Steering Angle ($64\times 64$) experiments, where it is compared with the DDPM and DDIM samplers. Stable Diffusion v1.5 (an advanced text-to-image model in our comparison) is fine-tuned on both RC-49 and UTKFace datasets, using the prompt templates “a photo of a chair at yaw angle [ANGLE] degrees” and “a portrait of the face of [AGE] year old”, respectively. The detailed implementation configurations are available in our GitHub repository.
Evaluation Setup. Consistent with the work of Ding et al. [ding2021ccgan, ding2023ccgan, ding2023efficient, ding2024turning], we predefine $m_{c}$ distinct points within the regression label range for each experiment, which we refer to as evaluation centers. For each evaluation center, we generate $N_{c}^{g}$ fake images, resulting in a total of 179,800, 60,000, 100,000, and 200,000 images generated by each candidate method in the RC-49, UTKFace, Steering Angle, and Cell-200 experiments, respectively. These images are evaluated using one overall metric and three separate metrics. The overall metric is the Sliding Fréchet Inception Distance (SFID) [ding2023ccgan], calculated as the average FID score across evaluation centers, with the standard deviation of these FIDs reported in Table  LABEL:tab:main_results. Additionally, we use Naturalness Image Quality Evaluator (NIQE) [mittal2012making] for visual fidelity, Diversity [ding2023ccgan] for image variety, and Label Score [ding2023ccgan] for label consistency. Note that the Diversity score cannot be computed for the Cell-200 dataset due to the lack of categorical image labels. For SFID, NIQE, and Label Score, lower values indicate better performance, while higher Diversity scores are preferred for greater variety in generated samples. For detailed evaluation setups, please refer to Appendix LABEL:sec:supp_detailed_test_setups.
IV-B Experimental Results
We evaluate candidate methods across seven experimental settings spanning four datasets and three resolutions, with complete quantitative results documented in Table LABEL:tab:main_results. Representative visual results are shown in Fig. LABEL:fig:line_graphs_rc49 and Fig. LABEL:fig:example_fake_images. Sampling cost comparisons for CCGM methods are performed on RC-49, with quantitative results detailed in Table LABEL:tab:ab_sampling_cost. The performance of DMD2-M against three other samplers is summarized in Table LABEL:tab:ab_effect_samplers. Our analysis of the experimental results highlights the following findings:
  • CCDM achieves the lowest SFID scores across all settings, indicating superior overall performance. Notably, in the RC-49 ($64\times 64$) experiment, its SFID score is half that of CcGAN and Dual-NDA, and only 5% of CcDPM’s, demonstrating significant image quality improvements.
  • Consistent with [ding2023ccgan, ding2024turning], class-conditional models perform significantly worse than CCGM. Notably, class-conditional GANs suffer from mode collapse in the Cell-200 experiment and frequent label inconsistency across settings.
  • Across all datasets except RC-49, CCDM achieves lower NIQE scores than the baseline CcGAN (SVDL+ILI), indicating superior visual fidelity. Even against Dual-NDA, designed to enhance CcGAN’s visual quality, CCDM outperforms in 4 of 7 settings.
  • The Diversity scores vary without a consistent trend. In some settings, CCDM achieves the highest diversity, while in others, it is notably lower yet remains above average.
  • In terms of label consistency, CCDM consistently outperforms or matches the performance of CcGAN and Dual-NDA across all settings.
  • In all settings, CCDM significantly outperforms CcDPM, and CcDPM generally lags behind GAN-based CCGM models. Notably, CcDPM shows poor visual quality and label consistency on RC-49, as evidenced by Fig. LABEL:fig:line_graphs_rc49, where CCDM outperforms it across all angles for evaluation in the $64\times 64$ experiment.
  • Table LABEL:tab:ab_effect_samplers shows that DDPM1K provides the best image quality but requires much longer sampling time, while DDIM250 offers the second-best results. Although DMD2-M sacrifices some image quality compared to DDIM250, it reduces sampling time from hours or days to seconds. Considering both Tables LABEL:tab:main_results and LABEL:tab:ab_effect_samplers, DMD2-M still outperforms CcGAN and Dual-NDA despite the quality trade-off.
  • While CCDM ($T^{\prime}=250$) achieves the lowest SFID scores, this comes at the cost of substantially higher computational requirements. Although DMD2-M demonstrates a viable trade-off between image quality and sampling efficiency (reducing both time and memory usage), developing more efficient sampling algorithms for CCDM remains an important direction for future research.